# Methods for Identifying Minimal Sufficient Statistics

### 11 Mar 2026 v1

@@ Abstract
Let $f_{\theta}$ be a joint density of a random sample $X$. A frequently used criterion asserts that a statistic $T(X)$ is minimal sufficient if, for any sample points $x$ and $y$, $T(x)=T(y)$ exactly when there exists a finite constant $h_{xy}>0$, independent of $\theta$, such that $f_{\theta}(y)=f_{\theta}(x)h_{xy}$ for every $\theta$. Although valid under regularity assumptions identified by Lehmann and Scheffé (1950), this statement is false as stated in general: we give a counterexample based on the choice of versions of Radon–Nikodym derivatives. We then propose a version-robust criterion that can be applied once sufficiency is established, and we use it to extend the method of Sato (1996) beyond Euclidean settings to analytic Borel sample spaces and standard Borel statistic spaces. While Lehmann–Scheffé–type regularity conditions and Sato’s approach can be difficult to verify in practice, our method is straightforward to check when the statistic is known to be sufficient. Finally, we analyze a distinct criterion for identifying minimal sufficient statistics developed by Pfanzagl (1994, 2017) and provide a discrete counterexample showing that it also requires additional assumptions.
Keywords: Analytic Borel spaces; minimal sufficiency; minimal sufficient statistics; Pfanzagl’s criterion; Sato’s method.
@@ 1 Introduction
According to the Lehmann–Scheffé theorem (Liese and Miescke, 2008, Theorem 7.17), one route to finding a uniformly minimum-variance unbiased estimator is to condition an unbiased estimator on a statistic that is both sufficient and complete. In practice, the main difficulty is that, apparently, there is no general procedure for constructing complete sufficient statistics. However, a useful observation is that, whenever a model admits a statistic that is sufficient and complete, every minimal sufficient statistic is automatically complete. Hence, methods for identifying minimal sufficient statistics provide an indirect but often practical route to complete sufficiency that is useful in applications such as estimation theory.
Let $\Theta$ be the parameter space and $T$ be a statistic defined on the statistical model. Informally, $T$ is said to be sufficient if, and only if, the data distribution conditional on $T$ does not depend on $\theta\in\Theta$ and $T$ is said to be minimal sufficient if, and only if, it is sufficient and, given any sufficient statistic $S$, there exists a measurable map $f$ such that $T=f\circ S$ almost everywhere. Throughout this paper, the formal definitions of statistical models, statistics, sufficiency and minimal sufficiency, standard and analytic Borel spaces, countably generated spaces and separable measurable spaces are collected in Appendix A.
In this paper, we discuss the limitations of two criteria for identifying minimal sufficiency and propose a method that avoids such limitations and also generalizes a method proposed by Sato (1996). The first and widely cited criterion considered in this paper can be found in texts such as Schervish (1995), see his Theorem 2.29, Wasserman (2004), see his Theorem 9.36, and Mavrakakis and Penzer (2021), see their Proposition 10.1.13, and it is stated as follows.
Criterion 1.1.
@@
In this formulation, the restriction applies over the entire (possibly uncountable) parameter space, which allows us to develop the Counterexample 2.1. Although widely quoted as a method for minimal sufficiency (see also Young and Smith, 2005; Makarov and Podkorytov, 2013; Olive, 2014; Gasperoni et al., 2025), Criterion 1.1 is often stated without explicit regularity hypotheses and is false in full generality. In particular, since densities are only defined almost everywhere, one can modify versions on null sets in a $\theta$-dependent way and thereby change the pointwise proportionality relation.
The formal conditions under which a reformulation of this method holds, preventing our counterexample, were first established by Sato (1996). This method is restricted to Euclidean spaces and is not simple to apply in practice. To bypass this practical difficulty, when a candidate statistic is already known to be sufficient (a condition that is typically easy to verify via the Neyman–Fisher factorization theorem), we introduce a more direct minimality method (Method 3.1). We then use this approach to extend Sato’s method beyond Euclidean spaces to more general settings, such as complete separable metric spaces (Method 3.2), and to establish a minimal-sufficiency method for exponential-family densities (Method 3.3).
@@
This criterion corresponds to Theorem 1.4.4 of Pfanzagl (1994). A later version of this result in Pfanzagl (2017) is stated with the additional requirement that $\{P_{\theta}\}_{\theta\in\Theta_{0}}$ is dense in $\{P_{\theta}\}_{\theta\in\Theta}$ w.r.t. the total variation distance, defined as $d(P,Q):=\sup_{E\in\Sigma}|P(E)-Q(E)|$. However, as noted at the beginning of the proof in Pfanzagl (1994), this density assumption is not necessary and may be omitted. Note also that Pfanzagl’s criterion imposes restriction only on at most a countable subset of the parameter space. However, it still does not hold without additional assumptions, as shown by our Counterexample 2.2.
The remainder of this paper is organized as follows. In Section 2, the limitations of these methods are demonstrated through specific counterexamples. Section 3 introduces the corrected and generalized methods that form the core contribution of this work, followed by examples illustrating their applications. Section 4 is dedicated to the formal proofs of these methods. Appendix A presents some basic definitions employed throughout the paper, namely, statistical models (Definition A.1), sufficient statistics (Definition A.2), standard Borel spaces (Definition A.3), countably generated spaces (Definition A.4), separable measurable spaces (Definition A.5), and analytic Borel spaces (Definition A.6). We denote the set of natural numbers by $\mathbb{N}=\{0,1,2,\ldots\}$. Moreover, we will abbreviate the expressions “if, and only if,” and “with respect to” by “iff” and “w.r.t.”, respectively.
@@ 2 Counterexamples
In this section, we construct counterexamples showing Criteria 1.1 and 1.2 require further regularity hypotheses. The first, similar to the example from Barndorff-Nielsen et al. (1976), perturbs the Gaussian density at a single $\theta$-dependent point, falsifying Criterion 1.1. The second invalidates Criterion 1.2 on a finite probability space.
@@
where $\mathbf{1}_{\mathbb{R}^{n}\setminus\{g(\theta)\}}$ denotes the indicator function of the set $\mathbb{R}^{n}\setminus\{g(\theta)\}$. Since $\lambda^{n}(\{g(\theta)\})=0$, $f_{\theta}$ is a joint density of $X:=(X_{1},\cdots,X_{n})$. Let $T:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be the identity function. Since $g$ is surjective, we infer that, for any $x,y\in\mathbb{R}^{n}$, we have $T(x)=T(y)$ iff there exists a finite constant $h_{xy}>0$ independent of $\theta$ such that $f_{\theta}(y)=f_{\theta}(x)h_{xy}$ for every $\theta\in\mathbb{R}$. (Indeed, if $x\neq y$, we can choose a $\theta\in\mathbb{R}$ such that $g(\theta)=y$ due to surjectivity. For this $\theta$, the indicator function in the definition of $f_{\theta}$ causes $f_{\theta}(y)=0$, while $f_{\theta}(x)\neq 0$, making the equality $f_{\theta}(y)=f_{\theta}(x)h_{xy}$ impossible for any $h_{xy}>0$.) Hence, according to Criterion 1.1, $X=T(X)$ is a minimal sufficient statistic, which implies, recalling that $\sum_{i=1}^{n}X_{i}$ is sufficient, that there exists a measurable function $u:\mathbb{R}\to\mathbb{R}^{n}$ such that $(x_{1},\cdots,x_{n})=u(\sum_{i=1}^{n}x_{i})$ $P_{\theta}$-a.e. for every $\theta\in\mathbb{R}$, where $P_{\theta}$ is the probability measure given by $P_{\theta}(E):=\int_{E}f_{\theta}d\lambda^{n}$. Since $f_{\theta}>0$ $\lambda^{n}$-a.e. for every $\theta\in\mathbb{R}$, we conclude that $(x_{1},\cdots,x_{n})=u(\sum_{i=1}^{n}x_{i})$ $\lambda^{n}$-a.e., a contradiction. To see this, define $A:=\{x\in\mathbb{R}^{n}:u(\sum_{i=1}^{n}x_{i})=x\}$ and $V:=\{x\in\mathbb{R}^{n}:\sum_{i=1}^{n}x_{i}=0\}$. It is straightforward to prove that $A$ is measurable and $(A-A)\cap(V\setminus\{0\})=\emptyset$, in which $A-A:=\{a_{1}-a_{2}:a_{1},a_{2}\in A\}$. However, if $\lambda^{n}(A)>0$, then, by Steinhaus Theorem, we know that AAA-A contains an open ball centered at 0, allowing us to conclude that $(A-A)\cap(V\setminus\{0\})\neq\emptyset$.
@@
We conclude this section with a simple observation. Since the counting measure on a non-empty finite set, normalized by its cardinality, is a probability measure, we may multiply the densities in Counterexample 2.2 by $4$; this shows that Criterion 1.2 remains false even if we additionally require $\mu$ to be a probability measure equivalent to $\{P_{\theta}\}_{\theta\in\Theta}$ (i.e., $\mu(E)=0$ if and only if $P_{\theta}(E)=0$ for every $\theta\in\Theta$).
@@ 3 Corrected and Generalized Methods
This section introduces correct criteria for identifying minimal sufficient statistics, together with examples illustrating their application. These criteria are rigorously proved in Section 4.
@@
Let $(\mathcal{X},\Sigma,\{P_{\theta}\}_{\theta\in\Theta})$ be a statistical model and $\mu:\Sigma\to\overline{\mathbb{R}}$ a $\sigma$-finite measure such that $P_{\theta}\ll\mu$ for every $\theta\in\Theta$. Suppose that $(\mathcal{X},\Sigma)$ is an analytic Borel space. For each $\theta\in\Theta$, let $f_{\theta}:\mathcal{X}\to\mathbb{R}$ be a density of $P_{\theta}$ w.r.t. $\mu$. Let $(\mathcal{T},\Sigma_{\mathcal{T}})$ be a standard Borel space and $T:\mathcal{X}\to\mathcal{T}$ a measurable function. For any $x\in\mathcal{X}$ and $\Theta^{\prime}\subseteq\Theta$, define
@@
Let $(\mathcal{X},\Sigma,\{P_{\theta}\}_{\theta\in\Theta})$ be a statistical model and $\mu:\Sigma\to\overline{\mathbb{R}}$ a $\sigma$-finite measure such that $P_{\theta}\ll\mu$ for every $\theta\in\Theta$. Suppose that $(\mathcal{X},\Sigma)$ is an analytic Borel space, $(\mathcal{T},\Sigma_{\mathcal{T}})$ is a standard Borel space, and $T:\mathcal{X}\to\mathcal{T}$ is a measurable function. For each $\theta\in\Theta$, let $f_{\theta}:\mathcal{X}\to\mathbb{R}$ be a density of $P_{\theta}$ w.r.t. $\mu$. Suppose that there exists a non-empty countable subset $\Theta_{0}\subseteq\Theta$ such that, for each $\theta\in\Theta$, there exists a sequence $(\theta_{n})_{n\in\mathbb{N}}$ of $\Theta_{0}$ such that the limit $\lim_{n\to\infty}f_{\theta_{n}}$ exists in $\mathbb{R}$ $\mu$-a.e. and for every $x\in\mathcal{X}$ we have
@@
For each $\theta>0$, define $\eta_{1}(\theta):=1/(k\theta)$ and $\eta_{2}(\theta):=-1/(2k\theta^{2})$. Let $a,b,c\in\mathbb{R}$ be arbitrary elements such that $a\eta_{1}(\theta)+b\eta_{2}(\theta)=c\,\,(i)$ for every $\theta>0$. Because of Method 3.3, to prove that $(\sum_{i=1}^{n}X_{i},\sum_{i=1}^{n}X_{i}^{2})$ is minimal sufficient, it suffices to prove that $a=b=c=0$. From $(i)$, we know that $2kc\theta^{2}-2a\theta+b=0$ for every $\theta>0$, allowing us to conclude, noting that this is a polynomial with degree at most $2$ in $\theta$, that 2kc=2a=b=02kc=-2a=b=0 and, hence, that $a=b=c=0$.

### 03 Apr 2026 v2

# Version-Robust Methods for Identifying Minimal Sufficient Statistics

@@ Abstract
Let $f_{\theta}$ be the joint density of a random sample $X$. A frequently used criterion asserts that a statistic $T(X)$ is minimal sufficient if, for any sample points $x$ and $y$, $T(x)=T(y)$ exactly when there exists a finite constant $h_{xy}>0$, independent of $\theta$, such that $f_{\theta}(y)=f_{\theta}(x)h_{xy}$ for all $\theta$. We show that this criterion is false in general via a counterexample exploiting the non-uniqueness of versions of Radon–Nikodym derivatives. Although Sato (1996) established sufficient regularity conditions for the validity of this criterion, these conditions are frequently intractable to verify in practice. We resolve this limitation by introducing a version-robust method applicable whenever sufficiency is known. Moreover, our method allows us to generalize Sato’s approach from Euclidean settings to arbitrary analytic Borel sample spaces and separable measurable statistic spaces. We also obtain a method for exponential-family densities under verifiable hypotheses. Taken together, these results clarify when pointwise likelihood-ratio arguments for minimal sufficiency are mathematically sound in irregular settings. Finally, we construct a counterexample demonstrating that a distinct criterion for minimal sufficiency due to Pfanzagl (1994, 2017) similarly fails in the absence of supplementary hypotheses. Identifying minimal sufficient statistics is important not only for parsimonious data reduction but also because, in models admitting complete sufficiency, such statistics provide a practical route to the complete sufficient structure underlying optimal estimation and prediction.
Keywords: Analytic Borel spaces; exponential families; irregular models; minimal sufficiency; minimal sufficient statistics; Pfanzagl’s criterion; Radon–Nikodym derivatives; Sato’s method.
@@ 1 Introduction
Minimal sufficiency is a data-reduction problem in dominated statistical models: one seeks a statistic that retains all inferentially relevant information in the sample while being minimal among sufficient reductions. A common strategy is to compare sample points through proportionality relations between likelihoods or densities. According to the Lehmann–Scheffé theorem (Liese and Miescke, 2008, Theorem 7.17), one route to finding a uniformly minimum-variance unbiased estimator is to condition an unbiased estimator on a statistic that is both sufficient and complete. In practice, however, complete sufficient statistics are often difficult to construct directly. Nevertheless, the relation between completeness and minimal sufficiency remains important: whenever a model admits a sufficient and complete statistic, every minimal sufficient statistic is automatically complete. For this reason, methods for identifying minimal sufficient statistics are useful not only for data reduction itself but also, in such models, as an indirect route to complete sufficiency in applications such as estimation theory. This connection is also relevant for prediction, since in such models an identified minimal sufficient statistic may serve as the complete sufficient statistic underlying the Bayes-optimal prediction procedures with frequentist coverage control developed in Hoff (2023).
Let $\Theta$ be the parameter space and $T$ be a statistic defined on the statistical model. Informally, $T$ is said to be sufficient if, and only if, the data distribution conditional on $T$ does not depend on $\theta\in\Theta$, and $T$ is said to be minimal sufficient if, and only if, it is sufficient and, given any sufficient statistic $S$, there exists a measurable map $f$ such that $T=f\circ S$ almost everywhere. Throughout this paper, the formal definitions of statistical models, statistics, sufficiency and minimal sufficiency, standard and analytic Borel spaces, countably generated spaces and separable measurable spaces are collected in Appendix A.
In this paper, we discuss the limitations of two pointwise criteria for identifying minimal sufficiency, propose version-robust methods that avoid such limitations, and also generalize a method proposed by Sato (1996). The first and widely cited criterion considered in this paper can be found in texts such as Schervish (1995), see his Theorem 2.29, Wasserman (2004), see his Theorem 9.36, and Mavrakakis and Penzer (2021), see their Proposition 10.1.13, and it is stated as follows.
Criterion $1.$
@@
In this formulation, the restriction applies over the entire (possibly uncountable) parameter space, which makes the pointwise proportionality relation sensitive to the chosen versions of the densities and allows us to develop Counterexample 2.1. Although widely quoted as a method for minimal sufficiency (see also Young and Smith, 2005; Makarov and Podkorytov, 2013; Olive, 2014; Gasperoni et al., 2025), Criterion 1.1 is often stated without explicit regularity hypotheses and is false in full generality. In particular, since densities are only defined almost everywhere, one can modify versions on null sets in a $\theta$-dependent way and thereby change the pointwise proportionality relation.
The formal conditions under which a reformulation of this method holds, preventing our counterexample, were first established by Sato (1996). This method is restricted to Euclidean spaces and is not simple to apply in practice. To bypass this practical difficulty, when a candidate statistic is already known to be sufficient (a condition that is typically easy to verify via the Neyman–Fisher factorization theorem), we introduce a more direct version-robust minimality method (Method 3.1). We then use this approach to extend Sato’s method beyond Euclidean spaces to analytic Borel sample spaces and separable measurable statistic spaces (Method 3.2), and to establish a minimal-sufficiency method for exponential-family densities (Method 3.3).
@@
This criterion corresponds to Theorem 1.4.4 of Pfanzagl (1994). A later version of this result in Pfanzagl (2017) is stated with the additional requirement that $\{P_{\theta}\}_{\theta\in\Theta_{0}}$ is dense in $\{P_{\theta}\}_{\theta\in\Theta}$ w.r.t. the total variation distance, defined as $d(P,Q):=\sup_{E\in\Sigma}|P(E)-Q(E)|$. However, as noted at the beginning of the proof in Pfanzagl (1994), this density assumption is not necessary and may be omitted. Note also that Pfanzagl’s criterion imposes a restriction only on at most a countable subset of the parameter space. However, it still does not hold without additional assumptions, as shown by our Counterexample $2.$
The remainder of this paper is organized as follows. In Section 2, the limitations of these methods are demonstrated through specific counterexamples that motivate the need for version-robust methods. Section 3 introduces the corrected and generalized methods that form the core contribution of this work, followed by examples illustrating their applications. Section 4 is dedicated to the formal proofs of these methods. Appendix A presents the measurable-space background used throughout the paper, namely, statistical models (Definition A.1), sufficient statistics (Definition A.2), standard Borel spaces (Definition A.3), countably generated spaces (Definition A.4), separable measurable spaces (Definition A.5), and analytic Borel spaces (Definition A.6). We denote the set of natural numbers by $\mathbb{N}=\{0,1,2,\ldots\}$. Moreover, we will abbreviate the expressions “if, and only if,” and “with respect to” by “iff” and “w.r.t.”, respectively.
@@ 2 Counterexamples
In this section, we construct counterexamples showing Criteria 1.1 and 1.2 require further regularity hypotheses. The first, similar to the example from Barndorff-Nielsen et al. (1976) and closely related to Example 1 of Taraldsen (2026), perturbs the Gaussian density at a single $\theta$-dependent point, falsifying Criterion $1.$ The second invalidates Criterion 1.2 on a finite probability space. These counterexamples motivate the methods introduced in Section 3.
@@
where $\mathbf{1}_{\mathbb{R}^{n}\setminus\{g(\theta)\}}$ denotes the indicator function of the set $\mathbb{R}^{n}\setminus\{g(\theta)\}$. Since $\lambda^{n}(\{g(\theta)\})=0$, $f_{\theta}$ is a joint density of $X:=(X_{1},\cdots,X_{n})$. Let $T:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be the identity function. Since $g$ is surjective, we infer that, for any $x,y\in\mathbb{R}^{n}$, we have $T(x)=T(y)$ iff there exists a finite constant $h_{xy}>0$ independent of $\theta$ such that $f_{\theta}(y)=f_{\theta}(x)h_{xy}$ for every $\theta\in\mathbb{R}$. (Indeed, if $x\neq y$, we can choose a $\theta\in\mathbb{R}$ such that $g(\theta)=y$ due to surjectivity. For this $\theta$, the indicator function in the definition of $f_{\theta}$ causes $f_{\theta}(y)=0$, while $f_{\theta}(x)\neq 0$, making the equality $f_{\theta}(y)=f_{\theta}(x)h_{xy}$ impossible for any $h_{xy}>0$.) Hence, according to Criterion 1.1, $X=T(X)$ is a minimal sufficient statistic, which implies, recalling that $\sum_{i=1}^{n}X_{i}$ is sufficient, that there exists a measurable function $u:\mathbb{R}\to\mathbb{R}^{n}$ such that $(x_{1},\cdots,x_{n})=u(\sum_{i=1}^{n}x_{i})$ $P_{\theta}$-a.e. for every $\theta\in\mathbb{R}$, where $P_{\theta}$ is the probability measure given by $P_{\theta}(E):=\int_{E}f_{\theta}d\lambda^{n}$. Since $f_{\theta}>0$ $\lambda^{n}$-a.e. for every $\theta\in\mathbb{R}$, we conclude that $(x_{1},\cdots,x_{n})=u(\sum_{i=1}^{n}x_{i})$ $\lambda^{n}$-a.e., a contradiction. To see this, define $A:=\{x\in\mathbb{R}^{n}:u(\sum_{i=1}^{n}x_{i})=x\}$ and $V:=\{x\in\mathbb{R}^{n}:\sum_{i=1}^{n}x_{i}=0\}$. It is straightforward to prove that AA is measurable and $(A-A)\cap(V\setminus\{0\})=\emptyset$, in which $A-A:=\{a_{1}-a_{2}:a_{1},a_{2}\in A\}$. However, if $\lambda^{n}(A)>0$, then, by Steinhaus Theorem, we know that AAA-A contains an open ball centered at 0, allowing us to conclude that $(A-A)\cap(V\setminus\{0\})\neq\emptyset$.
@@
We conclude this section with a simple observation. Since the counting measure on a non-empty finite set, normalized by its cardinality, is a probability measure, we may multiply the densities in Counterexample 2.2 by 44; this shows that Criterion 1.2 remains false even if we additionally require $\mu$ to be a probability measure equivalent to $\{P_{\theta}\}_{\theta\in\Theta}$ (i.e., $\mu(E)=0$ if and only if $P_{\theta}(E)=0$ for every $\theta\in\Theta$).
@@ 3 Corrected and Generalized Methods
This section introduces version-robust and correct criteria for identifying minimal sufficient statistics, together with examples illustrating their application. The main idea is to replace pointwise likelihood comparisons over the whole parameter space by arguments based on countable subfamilies or verifiable approximation hypotheses. These criteria are rigorously proved in Section 4.
These methods are particularly useful in settings in which pointwise likelihood comparisons are delicate but still informative, such as symmetry models, models with parameter-dependent support, and models on analytic Borel sample spaces. In particular, Examples 3.13.4 illustrate Method 3.1, Example 3.6 illustrates Method 3.2, and Example 3.7 illustrates Method $3.$
@@
Let $(\mathcal{X},\Sigma,\{P_{\theta}\}_{\theta\in\Theta})$ be a statistical model and $\mu:\Sigma\to\overline{\mathbb{R}}$ a $\sigma$-finite measure such that $P_{\theta}\ll\mu$ for every $\theta\in\Theta$. Suppose that $(\mathcal{X},\Sigma)$ is an analytic Borel space. For each $\theta\in\Theta$, let $f_{\theta}:\mathcal{X}\to\mathbb{R}$ be a density of $P_{\theta}$ w.r.t. $\mu$. Let $(\mathcal{T},\Sigma_{\mathcal{T}})$ be a separable measurable space and $T:\mathcal{X}\to\mathcal{T}$ a measurable function. For any $x\in\mathcal{X}$ and $\Theta^{\prime}\subseteq\Theta$, define
@@
Let $(\mathcal{X},\Sigma,\{P_{\theta}\}_{\theta\in\Theta})$ be a statistical model and $\mu:\Sigma\to\overline{\mathbb{R}}$ a $\sigma$-finite measure such that $P_{\theta}\ll\mu$ for every $\theta\in\Theta$. Suppose that $(\mathcal{X},\Sigma)$ is an analytic Borel space, $(\mathcal{T},\Sigma_{\mathcal{T}})$ is a separable measurable space, and $T:\mathcal{X}\to\mathcal{T}$ is a measurable function. For each $\theta\in\Theta$, let $f_{\theta}:\mathcal{X}\to\mathbb{R}$ be a density of $P_{\theta}$ w.r.t. $\mu$. Suppose that there exists a non-empty countable subset $\Theta_{0}\subseteq\Theta$ such that, for each $\theta\in\Theta$, there exists a sequence $(\theta_{n})_{n\in\mathbb{N}}$ of $\Theta_{0}$ such that the limit $\lim_{n\to\infty}f_{\theta_{n}}$ exists in $\mathbb{R}$ $\mu$-a.e. and for every $x\in\mathcal{X}$ we have
@@
For each $\theta>0$, define $\eta_{1}(\theta):=1/(k\theta)$ and $\eta_{2}(\theta):=-1/(2k\theta^{2})$. Let $a,b,c\in\mathbb{R}$ be arbitrary elements such that $a\eta_{1}(\theta)+b\eta_{2}(\theta)=c\,\,(i)$ for every $\theta>0$. Because of Method 3.3, to prove that $(\sum_{i=1}^{n}X_{i},\sum_{i=1}^{n}X_{i}^{2})$ is minimal sufficient, it suffices to prove that $a=b=c=0$. From $(i)$, we know that $2kc\theta^{2}-2a\theta+b=0$ for every $\theta>0$, allowing us to conclude, noting that this is a polynomial with degree at most 22 in $\theta$, that 2kc=2a=b=02kc=-2a=b=0 and, hence, that $a=b=c=0$.